OpenStudy (anonymous):

12m^2+8m factor completely

OpenStudy (anonymous):

Well did u get an answer yet?

OpenStudy (anonymous):

12m^2+8m 4m(3m+2) 4m=0 i.e m=0 3m+2=0 m=-2/3